# AffectGPT-R1: Leveraging Reinforcement Learning for Open-Vocabulary Multimodal Emotion Recognition

### v2

@@ Abstract
Open-Vocabulary Multimodal Emotion Recognition (OV-MER) aims to predict emotions without being constrained by predefined label spaces, enabling fine-grained emotion understanding. Unlike traditional discriminative methods, OV-MER leverages generative models, such as large language models, to capture the full spectrum of emotions and employs emotion wheels (EWs) for metric calculation. Previous approaches (e.g., AffectGPT) primarily rely on token-level loss during training. However, this objective is misaligned with the metrics used in OV-MER, while these metrics cannot be optimized via gradient backpropagation. In this paper, we propose AffectGPT-R1, a reinforcement learning framework that formulates EW-based metrics as a reward function and employs a policy-based optimization strategy to maximize this reward. Additionally, we introduce an extra reasoning process and investigate its necessity in OV-MER. To further refine model behavior, we incorporate auxiliary rewards that constrain both reasoning and emotion prediction. To prevent reward hacking, we propose to incorporate length penalties during training. Experimental results show that AffectGPT-R1 achieves substantial improvements on OV-MER. Beyond this task, our approach also enhances generalized emotion understanding, attaining state-of-the-art performance on MER-UniBench. To the best of our knowledge, this is the first work to adapt the R1-style methodology for emotion understanding, revealing the impact of reasoning processes and reinforcement learning in this domain. Our code is provided in the supplementary material and will be released to facilitate future research.
@@ 1 Introduction
Emotion is highly related to human cognition, decision-making, and behavior, playing a vital role in our daily life [picard2000affective, xie2024emovit]. Naturally, humans express emotions through multiple modalities, including (micro)-gestures, (micro)-expressions, audio tones, linguistic content, and other cues [ben2021video, chen2023smg, el2011survey, lian2024merbench]. To better understand human emotions, it is essential to integrate these clues, enabling models to capture emotional nuances. This has led to the development of Multimodal Emotion Recognition (MER) [lian2023mer, zhang2025moda]. Traditional MER approaches primarily rely on Ekman’s theory [ekman1992argument, li2023decoupled], which categorizes human emotions into six basic labels: anger, disgust, happiness, sadness, fear, and surprise. However, human emotions extend far beyond these basic categories, and such a limited label space inevitably results in inaccurate emotion descriptions [plutchik1980general, demszky2020goemotions].
@@
Recently, open-vocabulary MER (OV-MER) has emerged as a promising research direction, aiming to shift emotion recognition from constrained categories to the full spectrum of emotion, thereby enabling more nuanced emotion representations [lian2025ov, lian2025affectgpt]. To support this paradigm shift, OV-MER introduces new solutions and metrics. For solutions, it transitions from discriminative to generative approaches, leveraging the extensive vocabulary of large language models (LLMs) to expand the recognition scope of emotion labels. For metrics, it employs emotion wheels (EWs) to capture semantic relationships between distinct emotion words. Appendix A provides a detailed illustration of these metrics. To better address this task, current approaches primarily use token-level loss to align predicted and ground-truth labels [lian2025affectgpt, cheng2024emotion]. However, these methods suffer from a critical misalignment: token-level loss shows limited correlation with EW-based metrics. For instance, sadness and madness exhibit high token-level similarity but low semantic similarity in their underlying emotions. Furthermore, EW-based metrics cannot be directly optimized via gradient backpropagation, introducing significant challenges and limitations to existing solutions.
@@
In this paper, we propose AffectGPT-R1, which leverages EW-based metrics as the reward function and employs reinforcement learning to maximize this reward, enabling the direct optimization of these metrics. Meanwhile, we introduce an additional reasoning process and investigate its necessity in addressing OV-MER. To refine model behaviors, we incorporate auxiliary reward functions to constrain the model’s reasoning and emotion predictions. Through extensive experiments, we analyze the impact of different reward functions and identify the most effective ones. Additionally, during inference, the reward hacking problem [everitt2017reinforcement] arises, where the model tends to generate lengthy predictions with redundant emotional expressions (e.g., happy and happily). To mitigate this, we incorporate length penalties into the rewards, ensuring reward maximization while promoting more concise outputs. Experimental results demonstrate that AffectGPT-R1 achieves significant performance improvements on OV-MER. Our framework also attains state-of-the-art performance on MER-UniBench [lian2025affectgpt], a unified benchmark for generalized emotion understanding, further validating its effectiveness. To the best of our knowledge, this is the first work to apply R1-style training to OV-MER, demonstrating the effect of reasoning processes and reinforcement learning in this domain. The main contributions of this paper are summarized as follows:
@@
  • (Framework) We introduce AffectGPT-R1, the first work to apply an R1-style training strategy to OV-MER, shedding light on the role of reinforcement learning and reasoning processes in this task.
@@
  • (Reward) We systematically investigate various reward functions and conduct an in-depth analysis of their effectiveness, identifying the optimal rewards for OV-MER. To mitigate the reward hacking problem, we introduce length penalties, ensuring the model maximizes rewards while using the fewest emotional words.
@@
  • (Performance) We achieve state-of-the-art performance on OV-MER and MER-UniBench, proving the potential of reinforcement learning for emotion understanding.
@@ 2.1 Emotion Representation
Emotion representation defines the recognition scope of developed models and serves as the foundation of MER [gunes2011emotion]. There are two primary representation methods: dimensional emotion theory [schlosberg1954three] and basic emotion theory [ekman1992argument].
@@
This theory uses continuous dimensions to capture the core attributes underlying emotions. For example, Russell [russell1980circumplex] proposed a two-dimensional model, characterizing emotions based on valence (the pleasantness of an emotion) and arousal (the activation level of an emotion). In addition to this, some researchers introduced an additional dominance dimension to reflect the perceived control over emotions [russell1977evidence]. However, these dimensions are abstract and less intuitive for the general public, posing challenges for applications.
@@
This theory typically uses six or eight basic labels to describe emotional states. For example, Ekman [ekman1992argument] argued that human emotions can be categorized into six basic labels, while Plutchik [plutchik1980general] proposed eight primary emotions that serve as the foundation for all others. Nevertheless, human emotions are rich and diverse, extending far beyond these basic labels. Mapping such nuanced emotional expressions into a restricted set of basic labels inevitably leads to the loss of subtle emotional distinctions.
@@
To address these limitations, researchers propose an open-vocabulary representation manner [lian2025ov], which removes restrictions on the label space and enables models to predict any emotion. Unlike dimensional emotion theory, it focuses on more intuitive emotion words; Unlike basic emotion theory, it avoids restrictive label spaces, enabling more nuanced emotion representation. This paper focuses on open-vocabulary emotions and introduces AffectGPT-R1, a pioneering work that leverages the R1-style training strategy to tackle this task.
Figure 1: Overall pipeline of AffectGPT-R1. Our training pipeline consists of two phases: cold start training and reinforcement learning. In the first phase, we adopt the framework of AffectGPT but replace the generated content, including both thinking and answers. In the second phase, we propose five rewards and introduce length penalties to mitigate reward hacking, followed by GRPO for policy updates.
@@ 2.2 Reinforcement Learning
Reinforcement learning has advanced the upper-bound performance of LLMs, demonstrating impressive results in tasks such as mathematical verification and code execution [guo2025deepseek, shao2024deepseekmath, cobbe2021training]. This method primarily consists of two key components: policy optimization and reward modeling.
@@
During training, we aim to maximize the reward function through policy updates. Proximal Policy Optimization (PPO) [schulman2017proximal] is a widely used algorithm that leverages a value model for advantage estimation and incorporates token-wise KL penalties to prevent excessive optimization. Direct Preference Optimization (DPO) [rafailov2023direct] focuses on directly optimizing policies to align with human preferences, eliminating the need for an explicit reward model (as required in PPO) and instead relying on an implicit reward model. This approach reduces computational costs and engineering complexity. Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] is also a more efficient alternative to PPO, removing the value model and instead using a group-relative baseline computed from different outputs for the same input to guide optimization. Beyond these algorithms, some efforts have been made to enhance training efficiency and stability [hu2025reinforce++, yu2025dapo]. In this paper, we do not focus on designing novel policy optimization algorithms. Instead, we adopt GRPO for policy updates.
@@
Reward systems can be broadly categorized into rule-based and model-based rewards. The former is suitable for tasks with verifiable outcomes. For example, pass or fail unit tests in code generation [le2022coderl], output language requirements in question-answering tasks [guo2025deepseek], and IoU between predicted bounding boxes and ground-truth annotations in object detection [shen2025vlm]. The latter is mainly used for tasks where rewards are difficult to compute, such as safety, helpfulness, and harmlessness [zhang2024safetybench]. In prior work, how to design reward functions tailored for OV-MER remains unexplored. To this end, we introduce multiple rewards to supervise both thinking and answers, deriving insights into the effective rewards for OV-MER.
@@ 3 AffectGPT-R1
Our training pipeline consists of two phases: cold start training and reinforcement learning. In the first phase, we leverage a coarse-grained dataset to establish initial capabilities in emotion understanding and format alignment. In the second phase, we use a fine-grained dataset for reward calculation and policy updates. Figure 1 illustrates the overall training pipeline of AffectGPT-R1.
@@ 3.1 Notation Definition
This section introduces some necessary notations. Specifically, we use $v$ to represent the video, which contains multimodal clues (such as visual, acoustic, and lexical information), and $q$ to denote the user message. The output $o$ consists of two components: the thinking $o^{t}$ and the answer $o^{a}$. Specifically, $o^{t}$ and $o^{a}$ correspond to the content enclosed within the <<think>> and <</think>> tags and the <<answer>> and <</answer>> tags, respectively. We denote the ground-truth open-vocabulary labels as $y$. Figure 1 clarifies the meanings of these notations. In this paper, our main goal is to predict $y$ based on the video content $v$.
@@ 3.2 Cold Start Training
Human emotions are highly dependent on subtle cues [li2022deep, li2022cas], posing significant challenges for current models to capture all these signals and accurately interpret emotional states [lian2024gpt, lu2024gpt]. To enhance emotion understanding, we employ large-scale coarse-grained descriptive emotion datasets for cold-start training. In this phase, we adopt the same model architecture as AffectGPT [lian2025affectgpt] (see Figure 1). This framework leverages a pre-fusion layer to effectively integrate multimodal cues, emphasizing the inherently multimodal nature of human emotions. Unlike the original AffectGPT, which only outputs emotion words, we generate both thinking (enclosed within <<think>> and <</think>> tags) and answers (enclosed within <<answer>> and <</answer>> tags). Since descriptive emotion datasets already follow a thinking-like format [lian2023explainable], we treat these descriptions as the thinking content. Given our primary focus on OV-MER, we use open-vocabulary emotions as the answer content. Figure 1 illustrates an example of the input and output for cold-start training. This output format unifies descriptive and open-vocabulary emotions into a single task, enabling a single model to leverage their correlation and handle both tasks simultaneously. As demonstrated in Section 5.5, the quantity of cold-start data significantly affects both cold-start training performance and subsequent reinforcement learning. This further emphasizes our rationale for using large-scale datasets in this phase.
@@ 3.3.1 Policy Optimization
We employ GRPO [shao2024deepseekmath] for policy updates. Unlike PPO [schulman2017proximal], which relies on a critic model for advantage estimation, GRPO eliminates this component and instead uses group scores, significantly reducing memory overhead while improving training efficiency. Specifically, for a given input pair $(v,q)$, the policy model $\pi_{\theta_{\text{old}}}$ samples a group of responses $\mathbf{O}=\{o_{1},\cdots,o_{G}\}$, where $o_{i}$ is the $i$-th response. We then compute the reward for each response as $r_{i}=\mathcal{R}(o_{i}|v,q,y)$, where $\mathcal{R}(\cdot)$ is the reward function defined in Section 3.3.2. To evaluate the relative quality of $\{o_{i}\}_{i=1}^{G}$, we normalize the rewards by computing their mean and standard deviation, deriving the advantage as:
@@
During experiments, we observed significant divergence between the learned policy distribution and the initial model distribution, and the restrictive effect of the KL penalty term led to performance degradation. This trade-off between policy exploration and regularization has also been proven in prior work [everitt2017reinforcement]. To address this limitation, we remove the KL regularization term from our algorithm and adopt the following simplified objective function for policy update:
@@ 3.3.2 Reward Modeling
Reward functions guide the update direction of policy models. In this work, we propose five reward functions, all designed to evaluate thinking and answers. In Section 5.1, we examine the performance of the single reward and combinations of multiple rewards. Our goal is to identify the most effective reward strategy for OV-MER.
@@
We employ the official metrics of OV-MER to calculate accuracy rewards [lian2025ov]. Unlike discriminative MER, which operates within a fixed label space, OV-MER does not constrain the label space. Therefore, models are permitted to generate synonyms, i.e., emotions with the same meaning but expressed differently, which introduces challenges in evaluation. To mitigate the impact of synonyms, OV-MER introduces EW-based metrics, leveraging the structural information of emotion wheels for clustering. In these wheels, the inner section represents core emotions, while the outer section captures more nuanced emotions. Calculation details can be found in Appendix A:
@@
The accuracy reward ensures the correctness of $o^{a}$. However, without constraints on the thinking process, the generated output may contain unrelated $o^{a}$ and $o^{t}$. To this end, we introduce the alignment reward, ensuring that the reasoning process aligns with the final answer. Specifically, for each output, we extract emotion words $e^{t}$ from $o^{t}$ and compute their similarity with $o^{a}$:
@@
Similar to the alignment reward, this reward also imposes constraints on $o^{t}$. Specifically, we require the emotion words extracted from $o^{t}$ to match the ground-truth label $y$:
@@
Aforementioned rewards primarily evaluate text-level outputs $o$ but overlook whether the generated output aligns with the input video $v$. To assess perception correctness, inspired by EmoPrefer [lian2025emoprefer], we employ MLLMs to estimate pairwise preferences among outputs $(o_{i},o_{j})$ for a given video. Then, we apply the Bradley-Terry algorithm [hunter2004mm] to derive the ranking from these preferences. Finally, we assign a reward of 1 to outputs ranked in the top 50% and 0 to the rest. This reward is defined as follows:

### v3

@@ Abstract
Open-Vocabulary Multimodal Emotion Recognition (OV-MER) aims to predict emotions without being constrained by label spaces, enabling fine-grained emotion understanding. Unlike traditional discriminative methods, OV-MER leverages generative models to capture the full spectrum of emotions and employs emotion wheels (EWs) for metric calculation. Previous approaches (e.g., AffectGPT) primarily rely on token-level loss during training. However, this objective is misaligned with the metrics used in OV-MER, while these metrics cannot be optimized via gradient backpropagation. To address this limitation, we propose AffectGPT-R1, a reinforcement learning framework that treats EW-based metrics as a reward function and applies policy optimization to maximize this reward. Additionally, we introduce an explicit reasoning process and examine its necessity in OV-MER. To further guide model behavior, we incorporate auxiliary rewards that regularize both emotion reasoning and emotion prediction. We also apply length penalties to mitigate reward hacking. Experimental results demonstrate that AffectGPT-R1 yields significant performance improvements on OV-MER. Moreover, our approach enhances generalized emotion understanding, achieving state-of-the-art results on MER-UniBench. Our code is provided in the supplementary material and will be released to facilitate future research.
Machine Learning, ICML
@@ 1 Introduction
Emotion is highly related to human cognition, decision-making, and behavior, playing a vital role in our daily life (Picard, 2000; Xie et al., 2024). Naturally, humans express emotions through multiple modalities, including (micro)-gestures, (micro)-expressions, audio tones, linguistic content, and other cues (Ben et al., 2021; Chen et al., 2023a; El Ayadi et al., 2011; Lian et al., 2026b). To better understand human emotions, it is essential to integrate these clues, enabling models to capture emotional nuances. This has led to the development of Multimodal Emotion Recognition (MER) (Lian et al., 2023a; Zhang et al., 2025b). Traditional MER approaches primarily rely on Ekman’s theory (Ekman, 1992; Li et al., 2023), which categorizes human emotions into six basic labels: anger, disgust, happiness, sadness, fear, and surprise. However, human emotions extend far beyond these basic categories (Plutchik, 1980; Demszky et al., 2020), and such a restricted label space inevitably leads to imprecise emotion descriptions, thereby impairing the emotional intelligence of developed systems.
@@
Recently, open-vocabulary MER (OV-MER) has emerged as a promising research direction, aiming to shift emotion recognition from constrained categories to the full spectrum of emotion, thereby enabling more nuanced emotion representations (Lian et al., 2025a, c). To support this paradigm shift, OV-MER introduces new solutions and metrics. For solutions, it transitions from discriminative to generative approaches, leveraging the extensive vocabulary of large language models (LLMs) to expand the recognition scope of emotion labels. For metrics, it employs emotion wheels (EWs) to capture semantic relationships between distinct emotions. Appendix D provides a detailed illustration of these metrics. To better address this task, current approaches primarily use token-level loss to align predicted and true labels (Lian et al., 2025a; Cheng et al., 2024). However, these methods suffer from a critical misalignment: token-level loss shows limited correlation with EW-based metrics. For instance, sadness and madness exhibit high token-level similarity but low semantic similarity in their underlying emotions. Furthermore, EW-based metrics cannot be directly optimized via gradient backpropagation, introducing significant challenges and limitations to existing solutions.
@@
In this paper, we propose AffectGPT-R1, which leverages EW-based metrics as the reward function and employs reinforcement learning to maximize this reward, enabling the direct optimization of these metrics. Meanwhile, we introduce an additional reasoning process and investigate its necessity in addressing OV-MER. To refine model behaviors, we incorporate auxiliary reward functions to constrain the model’s reasoning and emotion predictions. Additionally, we observe the issue of reward hacking (Everitt et al., 2017), wherein the model tends to produce lengthy predictions containing redundant emotional expressions. To address this, we introduce length penalties into the reward function, thereby encouraging more concise outputs while still maximizing overall reward. Experimental results demonstrate that AffectGPT-R1 achieves significant improvements on OV-MER. Our framework also attains state-of-the-art performance on MER-UniBench (Lian et al., 2025a), a unified benchmark for generalized emotion understanding. Appendix B provides a more detailed discussion of our novelty. The main contributions of this paper are summarized below:
@@
  • (Framework) We introduce AffectGPT-R1, a pioneering work that applies reinforcement learning to OV-MER, shedding light on the role of reinforcement learning and reasoning mechanisms in this domain.
@@
  • (Reward) We systematically investigate various reward functions and conduct an in-depth analysis of their effectiveness. To mitigate reward hacking, we introduce length penalties, ensuring the model maximizes rewards while using the fewest emotional words.
@@
  • (Performance) We achieve state-of-the-art performance in open-vocabulary emotion detection and generalized emotion understanding, highlighting the potential of reinforcement learning in affective computing.
@@ 2.1 Emotion Representation
Emotion representation defines the recognition scope of developed models and serves as the foundation of MER (Gunes et al., 2011). There are two primary representation methods: dimensional emotion theory (Schlosberg, 1954) and basic emotion theory (Ekman, 1992).
@@
This theory uses continuous dimensions to capture the core attributes underlying emotions. For example, Russell (Russell, 1980) proposed a two-dimensional model, characterizing emotions based on valence (the pleasantness of an emotion) and arousal (the activation level of an emotion). In addition, some researchers have introduced a third dimension, dominance, to represent the degree of perceived control over emotions (Russell and Mehrabian, 1977). However, these dimensions are abstract and less intuitive for the general public, thus posing challenges in practical applications.
@@
This theory typically relies on six or eight basic labels to describe emotional states. For example, Ekman (Ekman, 1992) argued that human emotions can be categorized into six basic labels, whereas Plutchik (Plutchik, 1980) identified eight primary emotions that form the basis for other emotional experiences. Nevertheless, human emotions are rich and diverse, extending far beyond these basic labels. Mapping such nuanced emotional expressions into a restricted set of basic labels inevitably leads to the loss of subtle emotional distinctions.
@@
To address these limitations, researchers propose an open-vocabulary representation manner (Lian et al., 2025c), which removes restrictions on the label space and enables models to predict any emotion. Unlike dimensional emotion theory, it focuses on more intuitive emotion words; Unlike basic emotion theory, it avoids restrictive label spaces, enabling more nuanced emotion representation. Therefore, the open-vocabulary paradigm opens up new opportunities for affective computing. However, this field is still in its early stages, and effective solutions remain unclear. This paper presents a pioneering effort that leverages reinforcement learning to tackle this task.
Figure 1: Overall pipeline of AffectGPT-R1. Our training pipeline consists of two phases: cold start training and reinforcement learning. In the first phase, we adopt the framework of AffectGPT but replace the generated content, including both thinking and answers. In the second phase, we propose five rewards and introduce length penalties to mitigate reward hacking.
@@ 2.2 Reinforcement Learning
Reinforcement learning has advanced the upper-bound performance of LLMs, demonstrating impressive results in tasks such as mathematical verification and code execution (Guo et al., 2025; Shao et al., 2024; Cobbe et al., 2021). This method primarily consists of two key components: policy optimization and reward modeling.
@@
During training, we aim to maximize the reward function through policy updates. Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a widely used algorithm that leverages a value model for advantage estimation and incorporates token-wise KL penalties to prevent excessive optimization. Direct Preference Optimization (DPO) (Rafailov et al., 2023) focuses on directly optimizing policies to align with human preferences, eliminating the need for an explicit reward model and instead relying on an implicit reward model. This approach reduces computational costs and engineering complexity. Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a more efficient alternative to PPO, removing the value model and instead using a group-relative baseline computed from different outputs for the same input to guide optimization. Beyond these algorithms, some efforts have been made to enhance training efficiency and stability (Hu, 2025; Yu et al., 2025). This paper does not focus on designing new policy optimization algorithms, but rather on crafting optimal rewards for emotion recognition.
@@
Reward systems can be broadly categorized into rule-based and model-based rewards. The former is suitable for tasks with verifiable outcomes. For example, pass or fail unit tests in code generation (Le et al., 2022), output language requirements in question-answering tasks (Guo et al., 2025), and IoU between predicted bounding boxes and ground-truth annotations in object detection (Shen et al., 2025). The latter is mainly used for tasks where rewards are difficult to compute, such as safety, helpfulness, and harmlessness (Zhang et al., 2024). In prior work, how to design reward functions tailored for OV-MER remains unexplored. To this end, we introduce multiple rewards to supervise both thinking and answers, deriving insights into the effective rewards for this task.
@@ 3 AffectGPT-R1
Our training pipeline consists of two phases: cold start training and reinforcement learning. In the first phase, we utilize a large-scale, coarse-grained dataset to establish initial capabilities in emotion understanding and format alignment. In the second phase, we use a high-quality, fine-grained dataset for reward calculation and policy updates. Figure 1 illustrates the overall pipeline of AffectGPT-R1.
@@ 3.1 Notation Definition
This section introduces some necessary notations. Specifically, we use $v$ to represent the video, which contains multimodal clues (such as visual, acoustic, and lexical information), and $q$ to denote the user message. The output $o$ consists of two components: the thinking $o^{t}$ and the answer $o^{a}$. Specifically, $o^{t}$ and $o^{a}$ correspond to the content enclosed within the <<think>> and <</think>> tags and the <<answer>> and <</answer>> tags, respectively. We denote the ground-truth open-vocabulary labels as $y$. Figure 1 clarifies the meanings of these notations. Our main goal is to predict the emotion $y$ based on the video content $v$. We would like to emphasize that this paper focuses on utterance-level emotion recognition, ensuring fair comparison with prior work (Lian et al., 2025a, c). Appendix G.4 provides additional clarification regarding our task formulation.
@@ 3.2 Cold Start Training
Human emotions are highly dependent on subtle cues (Li et al., 2022a, b), posing significant challenges for current models to capture all these signals and accurately interpret emotional states (Lian et al., 2024b; Lu et al., 2024). To enhance emotion understanding, we employ large-scale coarse-grained descriptive emotion datasets for cold-start training. In this phase, we adopt the same model architecture as AffectGPT (Lian et al., 2025a) (see Figure 1). This framework leverages a pre-fusion layer to effectively integrate multimodal cues, emphasizing the inherently multimodal nature of human emotions. Unlike the original AffectGPT, which only outputs emotion words, we generate both thinking (enclosed within <<think>> and <</think>> tags) and answers (enclosed within <<answer>> and <</answer>> tags). Since descriptive emotion datasets already follow a thinking-like format (Lian et al., 2023b), we treat these descriptions as the thinking content. Given our primary focus on OV-MER, we use open-vocabulary emotions as the answer content. Figure 1 illustrates an example of the input and output for cold-start training. This output format unifies descriptive and open-vocabulary emotions into a single task, enabling a single model to leverage their correlation and handle both tasks simultaneously. As demonstrated in Section 5.5, the quantity of cold-start data significantly affects both cold-start training performance and subsequent reinforcement learning. This further emphasizes our rationale for using large-scale datasets in this phase.
@@ 3.3.1 Policy Optimization
We employ GRPO (Shao et al., 2024) for policy updates. Unlike PPO (Schulman et al., 2017), which relies on a critic model for advantage estimation, GRPO eliminates this component and instead uses group scores, significantly reducing memory overhead while improving training efficiency. Specifically, for a given input pair $(v,q)$, the policy model $\pi_{\theta_{\text{old}}}$ samples a group of responses $\mathbf{O}=\{o_{i}\}_{i=1}^{G}$. We then compute the reward for each response as $r_{i}=\mathcal{R}(o_{i}|v,q,y)$, where $\mathcal{R}(\cdot)$ is the reward function defined in Section 3.3.2. To evaluate their relative quality, we normalize the rewards by computing their mean and standard deviation:
@@
During experiments, we observed a divergence between the learned policy distribution and the initial model distribution, and the restrictive nature of the KL penalty term constrained overall performance. This trade-off between policy exploration and regularization has been proven in prior work (Yu et al., 2025). Following the prior work (Yu et al., 2025), we remove the KL regularization term and adopt the simplified objective function for policy update. Experimental results in Section 5 demonstrate the effectiveness of our method.
@@ 3.3.2 Reward Modeling
Reward functions guide the update direction. This paper proposes five reward functions, all designed to steer thinking and answers. Section 5.1 reveals the performance of the single reward and combinations of multiple rewards.
@@
We employ the official metrics of OV-MER to compute accuracy rewards (Lian et al., 2025c). Unlike discriminative MER, which operates within a fixed label space, OV-MER imposes no constraints on the label space. Consequently, models are allowed to generate synonyms, i.e., emotions with identical meanings but different expressions, which poses challenges for evaluation. To mitigate these challenges, OV-MER introduces EW-based metrics, which leverage the structural information of emotion wheels for clustering. Following prior work (Lian et al., 2025c), we utilize five emotion wheels (see Appendix C), and a detailed calculation procedure for EW-based metrics is presented in Appendix D. We denote this reward as:
@@
Accuracy reward ensures the correctness of $o^{a}$. However, without constraints on the reasoning process, the generated output may include irrelevant $o^{a}$ and $o^{t}$. To address this, we introduce alignment reward, which ensures that the reasoning process is consistent with the final answer. Specifically, we extract emotion words $e^{t}$ from $o^{t}$ and compute their similarity with $o^{a}$. Appendix E provides the detailed methods for emotion word extraction and similarity calculation. We denote this reward as:
@@
Similar to alignment reward, this reward also imposes constraints on $o^{t}$. Specifically, we require the emotion words extracted from $o^{t}$ to match the true label $y$:
@@
Aforementioned rewards primarily evaluate text-level outputs $o$ but overlook whether the generated output aligns with the input video $v$. To assess perception correctness, inspired by EmoPrefer (Lian et al., 2026a), we employ MLLMs to estimate pairwise preferences among outputs $(o_{i},o_{j})$ for a given video. Then, we apply the Bradley-Terry algorithm (Hunter, 2004) to derive the ranking from these preferences. Finally, we assign a reward of 1 to outputs ranked in the top 50% and 0 to the rest. This reward is defined as follows: